OpenStudy (anonymous):

solve for x? 3(2x+2) - 4 = 5(x+1)

OpenStudy (anonymous):

3(2x +3) - 4 = 5(x+1) 6x +6 -4 = 5x+ 5 6x +2 = 5x + 5 6x - 5x = 5 - 2

OpenStudy (lauluisa):

6x + 6 -4 = 5x + 5 Then you add like terms.. 6x +2 = 5x + 5 -5x -5x x + 2 = 5 - 2 -2 You substract 2, because you have to move it to the other side, so it becomes a negative. x = 3 Hope it helps c: